### v1

@@ 2.3 Models
For the model al03K, we adopt the spatially-uniform heating rate described by a four-segment broken power-law function of time given by Equation 3 of Klion et al. (2022). Compared to our fiducial heating rate implementation, we will show in Section 6 that this prescription may have overestimated (underestimated) heating in early (late) phase and will discuss the consequence on the resulting impact of heating.
@@
Matter from the grid or tracer particles are considered unbound when their Bernoulli parameter satisfies $B_{e}>0$ and reach radial distances $\geq 10^{8}$ cm. The Bernoulli parameter is defined as:
@@
Table 1: Models investigated. The first five columns show the model name and their set-up parameters; from left to right are the model names, viscosity parameter $\alpha$, whether r heating is included, threshold temperature for heating $T_{\rm heat}$, and the resolution for interpolation in $Y_{e,T6}$. The following three columns are the resulting unbound mass $M_{ej}$, mass averaged radial velocity at $T=1$ GK $\langle v_{r}\rangle_{T9}$ and $r$-process heating energy $Q_{r}$. See Section 2.3 for the definition of unbound. The last two columns are the ratio of the mass averaged radial velocity at $T=1$ GK $\langle v_{r}\rangle_{T9}$ compared to the corresponding values in the no-heating (nH) models, we separate the tracers in two groups, $\mathcal{R}_{\langle v_{r}\rangle;<0.25}$ for low-$Y_{e,T6}$ tracers, and $\mathcal{R}_{\langle v_{r}\rangle;\geq 0.25}$ for high-$Y_{e,T6}$ tracers.
@@ 3 Overall evolution
The accretion in the inner region and the expansion at the outer region lead to a decrease in density, which in turn reduces the efficiency of neutrino cooling on the viscous timescale ($\sim 300$ ms for $\alpha=0.03$). As neutrino cooling becomes insufficient to offset viscous heating, gas pressure once again supports the inner disk against gravity, thus slowing down the accretion rate. At the same time, the remaining material in the outer region gains energy from viscous heating and nuclear recombination, thus accelerating its expansion. The launching of this outflow occurs within approximately $t\leq 1$ s and results in a decrease in both temperature and density of the outflowing matter (Fernández & Metzger, 2013). As the internal energy of the outermost material is gradually converted into kinetic energy, the kinetic energy rises again between 0.2 and 1 s and becomes the dominant component of the system’s energy budget. This transition is presented in the left panel of Figure 3, which depicts the energy evolution across seven models. Solid lines represent the fluid’s total kinetic energy, while dashed lines represent the total internal energy. The different line colors denote various models, as indicated in the legend.
Figure 2 shows that the model which includes nuclear heating launches the outflow on a shorter timescale. By $t=3.6$ s, only a small amount of mass remains in model al03T4. In contrast, the model without heating (al03nH) retains a significant amount of mass, indicating a slower and less efficient outflow. The model with $r$-process heating also produces a more spherical outflow, particularly evident at $t=1.6$ s. This more spherical morphology suggests that the radial velocity component dominates over others, corresponding to less convection and a more rapid expansion.
@@
4 Tracer evolution and $r$-process heating feedback
Figure 4: The upper four panels depict the distribution of tracers from 100 ms to 16 s (left to right). The color map represents each tracer’s $Y_{e,T6}$ value. The empty region in the left panel (at t = 0.1 s) corresponds to the initial torus, where temperature exceeds 6 GK. In this high-temperature region, tracers are not yet assigned with their $Y_{e,T6}$ value. Heating rate is also not required under such conditions. The lower panels display the specific nuclear heating rate derived from tracers corresponding to the upper panels. Heating regions (outflows) are shown in red, while cooling regions (inflows) are in blue.
To better understand the heating implementation, we track the evolution of passive tracers in our simulations, which are initially distributed according to the density profile, and analyze their behavior. Initially, tracers originating from the torus outer regions rapidly reach the critical temperature threshold of $T=6$ GK while maintaining its low $Y_{e}$. As shown in the upper panels of Figure 4, the value of $Y_{e,T6}$ of these early, fast-moving tracers remains low. This condition leads to high nuclear heating rates (Figure 1) and an earlier onset of heating contributions. The lower panels of Figure 4 show that a strong heating rate (indicated in red) first appears in those regions corresponding to the locations of low-$Y_{e}$ tracers identified in the upper panel. These low-$Y_{e}$ tracers contribute to the early energy input into the system and are efficiently accelerated outward.
@@ 4 Tracer evolution and $r$-process heating feedback
Despite the local variations in heating rates associated with the tracers’ composition, the global morphology of the ejecta remains nearly spherical. This spherical structure is caused by the convective and turbulent motion, which efficiently redistributes tracers in different latitudes. As a consequence, the large-scale heating rate distribution smoothes out over time, and the energy is also mixed throughout the outflow, keeping the total ejecta shape spherical. In Figure 4, the upper panels show the spatial mixing of tracer components, and the lower panels show how the heating distribution becomes isotropic over time.
@@
5 Interplay of viscous heating and $r$-process heating
@@
During the initial phase of the simulation (within hundreds of milliseconds), both internal and kinetic energy decrease due to accretion, as described in Section 3. Models with $\alpha=0.06$ evolve more rapidly than those with $\alpha=0.03$, as reflected in the earlier decline of both internal and kinetic energies. This faster evolution is a consequence of a more efficient energy transport associated with higher viscosity parameter. High $\alpha$ also causes faster conversion of thermal energy into kinetic energy and an earlier rise in the kinetic energy at around $t=0.2$ s.
In the early stage of evolution (before $5\times 10^{-3}$ s), the heating rate for model al03T6 is approximately six times greater than that of al03T4 and al06T4. This difference is due to the heating onset threshold being set at a temperature of 6 GK rather than 4 GK, enabling earlier heating in model al03T6. Across all models, the total heating rate becomes negative between $5\times 10^{-3}$ and $4\times 10^{-2}$ s due to notable inward matter flow, which induces cooling as described in Section 2.2. Subsequently, the total heating rate becomes positive again and peaks between 4 and 10 s depending on the model, followed by a decline consistent with the specific heating rate profiles presented in Figure 1.
@@
The radial velocity distributions at $t=0.6$ s indicate maximum ejecta velocities of approximately 0.1 c, with most models peaking around 0.02 c and an extended tail toward lower velocities. Models with $\alpha=0.06$ exhibit faster ejecta, the peak velocities are $\sim 0.02$ c and $\sim 0.1$ c for models without and with heating, respectively. These models expand more rapidly into the outer, lower-temperature regions. The upper panels of Figure 7 illustrate that viscosity dominates over $r$-process heating during the early phase (within the first few hundred milliseconds).
@@
By t = 6 s, the entropy distribution in models without nuclear heating (al03nH and al06nH) shows only minor changes relative to that at $t=0.6$ s. In contrast, models that include $r$-process heating display a five-fold increase in entropy, caused by the additional energy input from nuclear heating. The radial velocity distributions indicate an overall increase, peaking at approximately 0.15 c, highlighting the acceleration of the ejecta due to $r$-process heating feedback. Models al06T4 and al03T6 even achieve maximum radial velocities of 0.25 c. Velocity distributions of models without heating only slightly increase relative to $t=0.6$ s due to thermal energy conversion into kinetic energy. Over time, the heating rate deposits energy into the fluid, and on a timescale of seconds, heating alters the thermodynamic properties of the ejecta. The results in Figure 7 suggest that while viscous heating governs the early dynamical evolution, $r$-process heating becomes the dominant factor after a few seconds.
@@
Figure 8: Tangential velocity of unbound tracers as a function of radius at t = 6 s. Tracers from models with and without nuclear heating are shown in different colors, as in the legend. The upper and lower panels correspond to two viscosity parameters, $\alpha=0.03$ and $\alpha=0.06$, respectively..
@@
Figure 8 shows the tangential velocity $v_{t}$ of unbound tracers at t = 6 s for four models. In the inner region, $r\lesssim 10^{5}$ km, tracers from models without nuclear heating (al03nH and al06nH) exhibit significantly higher $v_{t}$. In contrast, for models that include heating, $v_{t}$ remains nearly flat within this region, indicating that convection is suppressed. This behavior is consistent with the low-velocity end of the radial velocity histograms in the lower panel of Figure 7, where the models without heating contain a greater amount of low-velocity ejecta than the heating models at t = 6 s. These slow-moving ejecta reflect convective motion near the BH.
@@
Figure 9: Time evolution of the unbound mass reaching $r=10^{8}$ cm, as defined in Equation (1), for various models. Each model is represented by a different color, as indicated in the legend.
@@
We analyze tracers that are initially distributed with the fluid density. Following the gravitational binding criterion defined in Section 2.3, Figure 9 shows the temporal evolution of the cumulative unbound mass reaching $r=10^{8}$ cm for different models, as indicated in the legend. Models that include heating exhibit rapid unbound mass growth, reaching saturation within $\sim 0.5$ s. In contrast, non-heating (nH) models gradually increase over roughly 2 s. This points out that neglecting $r$-process heating results in slower ejection of material, with the saturation timescale differing by 1.5 s. The final unbound mass values ($M_{ej}$) are listed in Table 1, showing a 10% increase in models that include heating. Additionally, models with higher viscosity ($\alpha=0.06$) evolve more quickly and eject 10% more mass than those with $\alpha=0.03$, consistent with the expectation that enhanced angular momentum transport facilitates earlier mass ejection.
6 Comparison with time-dependent uniform heating rate
@@ 7 DISCUSSION
Angular momentum transport in realistic merger disks is expected to occur via magnetohydrodynamic (MHD) turbulence driven by the magnetorotational instability (e.g., Balbus & Hawley 1998) instead of shear viscosity. Previous long-term 3D MHD simulations of BH accretion disks from NS mergers have shown that such turbulence combined with large-scale magnetic stresses can drive faster and more massive disk winds for strong poloidal fields, with typical average outflow velocities around 0.1 c instead of 0.03 c for viscous hydrodynamic models (Siegel & Metzger, 2018; Fernández et al., 2019; Miller et al., 2019; Just et al., 2022; Hayashi et al., 2022). In this regime of higher baseline velocities, the relative impact of r-process heating is likely to be reduced, leading to more moderate velocity increases of the order of $\sim 10\%$, if the energy gain per baryon remains constant, compared to the factor of two enhancement observed in our simulations. For toroidal fields or weak poloidal fields, average outflow velocities are closer to the viscous hydrodynamic values (Christie et al., 2019; Fahlman & Fernández, 2022; Hayashi et al., 2023), in which case $r$-process heating can still be an order unity correction to the outflow velocities. The field geometry immediately after the merger is complex, with a mixture of poloidal and toroidal components spanning a range of spatial scales due to turbulence (e.g., Most et al. 2021; Aguilera-Miret et al. 2023; Izquierdo et al. 2024), hence the definitive impact of r-process heating on outflow velocities with realistic field geometries remains an area of active study.
Our current heating prescription interpolates the nuclear energy release based on $Y_{e,T6}$ and the instantaneous temperature. However, outflows from merger disks exhibit a broad range of entropies and expansion timescales, which also affect nucleosynthesis and heating rates. Expanding the heating tables to include these additional parameters in future work can improve the physical accuracy of the heating rate.

### v2

@@ II.3 Models
For the model al03K, we adopt the spatially-uniform heating rate described by a four-segment broken power-law function of time given by Equation 3 of Klion et al. (2022). Compared to our fiducial heating rate implementation, we will show in Section VI that this prescription may have overestimated (underestimated) heating in early (late) phase and will discuss the consequence on the resulting impact of heating.
@@
Matter from the grid or tracer particles are considered unbound when their Bernoulli parameter satisfies $B_{e}>0$ and reach radial distances $\geq 10^{3}$ km. The Bernoulli parameter is defined as:
@@
Table 1: Models investigated. The first five columns show the model name and their set-up parameters; from left to right are the model names, viscosity parameter $\alpha$, whether r heating is included, threshold temperature for heating $T_{\rm heat}$, and the resolution for interpolation in $Y_{e,T6}$. The following three columns are the resulting unbound mass $M_{ej}$, mass averaged radial velocity at $T=1$ GK $\langle v_{r}\rangle_{T9}$ and $r$-process heating energy $Q_{r}$. See Section II.3 for the definition of unbound. The last two columns are the ratio of the mass averaged radial velocity at $T=1$ GK $\langle v_{r}\rangle_{T9}$ compared to the corresponding values in the no-heating (nH) models, we separate the tracers in two groups, $\mathcal{R}_{\langle v_{r}\rangle;<0.25}$ for low-$Y_{e,T6}$ tracers, and $\mathcal{R}_{\langle v_{r}\rangle;\geq 0.25}$ for high-$Y_{e,T6}$ tracers.
@@ III Overall evolution
The accretion in the inner region and the expansion at the outer region lead to a decrease in density, which in turn reduces the efficiency of neutrino cooling on the viscous timescale ($\sim 300$ ms for $\alpha=0.03$). As neutrino cooling becomes insufficient to offset viscous heating, gas pressure once again supports the inner disk against gravity, thus slowing down the accretion rate. At the same time, the remaining material in the outer region gains energy from viscous heating and nuclear recombination, thus accelerating its expansion. The launching of this outflow occurs within approximately $t\leq 1$ s and results in a decrease in both temperature and density of the outflowing matter (Fernández and Metzger, 2013). As the internal energy of the outermost material is gradually converted into kinetic energy, the kinetic energy rises again between 0.2 and 1 s and becomes the dominant component of the system’s energy budget. This transition is presented in the left panel of Figure 3, which depicts the energy evolution across seven models. Solid lines represent the fluid’s total kinetic energy, while dashed lines represent the total internal energy. The different line colors denote various models, as indicated in the legend.
Figure 2 shows that the model which includes nuclear heating launches the outflow on a shorter timescale. By $t=3.6$ s, only a small amount of mass remains in model al03T4. In contrast, the model without heating (al03nH) retains a significant amount of mass, indicating a slower and less efficient outflow. The model with $r$-process heating also produces a more spherical outflow, particularly evident at $t=1.6$ s. The global increase in thermal energy drives this spherical morphology through rapid, isotropic expansion, which smooths large scale asymmetries and makes radial velocity dominant over tangential convective fluctuation.
@@
IV Tracer evolution and $r$-process heating feedback
Figure 4: The upper four panels depict the distribution of tracers from 100 ms to 16 s (left to right) for model al03T4. The color map represents each tracer’s $Y_{e,T6}$ value. The empty region in the left panel (at t = 0.1 s) corresponds to the initial torus, where temperature exceeds 6 GK. In this high-temperature region, tracers are not yet assigned with their $Y_{e,T6}$ value. Heating rate is also not required under such conditions. The lower panels display the specific nuclear heating rate derived from tracers corresponding to the upper panels. Heating regions (outflows) are shown in red, while cooling regions (inflows) are in blue.
To better understand the heating implementation, we track the evolution of passive tracers in our simulations, which are initially distributed according to the density profile, and analyze their behavior. Initially, tracers originating from the torus outer regions rapidly reach the critical temperature threshold of $T=6$ GK while maintaining its low $Y_{e}$. As shown in the upper panels of Figure 4, the value of $Y_{e,T6}$ of these early, fast-moving tracers in model al03T4 remains low. This condition leads to high nuclear heating rates (Figure 1) and an earlier onset of heating contributions. The lower panels of Figure 4 show that a strong heating rate (indicated in red) first appears in those regions corresponding to the locations of low-$Y_{e}$ tracers identified in the upper panel. These low-$Y_{e}$ tracers contribute to the early energy input into the system and are efficiently accelerated outward.
@@ IV Tracer evolution and $r$-process heating feedback
Despite the local variations in heating rates associated with the tracers’ composition, the global morphology of the ejecta remains nearly spherical. Convection and turbulence redistribute tracers across different latitudes, producing spatially uniform heating. This uniform heating deposited into the ejecta, increases the thermal pressure and drives rapid expansion, thereby shaping the spherical ejecta morphology. In Figure 4, the upper panels show the spatial mixing of tracer components, and the lower panels show how the heating distribution becomes isotropic over time.
@@
V Interplay of viscous heating and $r$-process heating
@@
During the initial phase of the simulation (within hundreds of milliseconds), both internal and kinetic energy decrease due to accretion, as described in Section III. Models with $\alpha=0.06$ evolve more rapidly than those with $\alpha=0.03$, as reflected in the earlier decline of both internal and kinetic energies. This faster evolution is a consequence of a more efficient energy transport associated with higher viscosity parameter. High $\alpha$ also causes faster conversion of thermal energy into kinetic energy and an earlier rise in the kinetic energy at around $t=0.2$ s.
In the early stage of evolution (before $5\times 10^{-3}$ s), the heating rate for model al03T6 (green line) is approximately six times greater than that of al03T4 and al06T4 (blue and red lines). This difference is due to the heating onset threshold being set at a temperature of 6 GK rather than 4 GK, enabling earlier heating in model al03T6. Across all models, the total heating rate becomes negative between $5\times 10^{-3}$ and $4\times 10^{-2}$ s due to notable inward matter flow, which induces cooling as described in Section II.2. Subsequently, the total heating rate becomes positive again and peaks between 4 and 10 s depending on the model, followed by a decline consistent with the specific heating rate profiles presented in Figure 1.
@@
The radial velocity distributions at $t=0.6$ s indicate maximum ejecta velocities of approximately 0.1 c, with most models peaking around 0.02 c and an extended tail toward lower velocities. Models with $\alpha=0.06$ exhibit faster ejecta, the peak velocities are $\sim 0.02$ c and $\sim 0.1$ c for models without and with heating (light red filled bars and red outlined bars), respectively. These models expand more rapidly into the outer, lower-temperature regions. The upper panels of Figure 7 illustrate that viscosity dominates over $r$-process heating during the early phase (within the first few hundred milliseconds).
@@
By t = 6 s, the entropy distribution in models without nuclear heating (al03nH and al06nH, blue and red filled bars) shows only minor changes relative to that at $t=0.6$ s. In contrast, models that include $r$-process heating display a five-fold increase in entropy, caused by the additional energy input from nuclear heating. The radial velocity distributions indicate an overall increase, peaking at approximately 0.15 c, highlighting the acceleration of the ejecta due to $r$-process heating feedback. Models al06T4 and al03T6 (red and green outlined bars) even achieve maximum radial velocities of 0.25 c. Velocity distributions of models without heating only slightly increase relative to $t=0.6$ s due to thermal energy conversion into kinetic energy. Over time, the heating rate deposits energy into the fluid, and on a timescale of seconds, heating alters the thermodynamic properties of the ejecta. The results in Figure 7 suggest that while viscous heating governs the early dynamical evolution, $r$-process heating becomes the dominant factor after a few seconds.
@@
Figure 8: Ratio of the tangential velocity to the angle-averaged radial velocity for unbound tracers as a function of radius. Tracers from models with and without nuclear heating are shown in different colors, as indicated in the legend. The upper and lower panels correspond to two viscosity parameters, $\alpha=0.03$ and $\alpha=0.06$, respectively. Snapshots are selected for each model at times corresponding to similar positions of the forward shock front, ensuring a consistent comparison of the outflow structure.
@@
Figure 8 shows the ratio of tangential velocity to angle-averaged radial velocity $v_{t}/\langle v_{r}\rangle$ of unbound tracers at t = 6 s for models without heating (al03nH and al06nH), t = 2.4 s for model al03T4, and t = 2.8 s for model al06T4. In the inner region, $r\lesssim 10^{5}$ km, tracers from models without nuclear heating (al03nH and al06nH, green and orange dots) exhibit significant dispersion in $v_{t}/\langle v_{r}\rangle$ around zero, indicating strong convective activity. In contrast, for models that include heating (blue and red dots), $v_{t}/\langle v_{r}\rangle$ remains closer to zero in this region, indicating that convection is suppressed. However, this suppression primarily affects convection near the marginally-bound ejecta close to the black hole, and convective and turbulent motions remain active on larger scales, driving the overall spherical morphology of the ejecta. This behavior is consistent with the low-velocity end of the radial velocity histograms in the lower panel of Figure 7, where the models without heating contain a greater amount of low-velocity ejecta than the heating models at t = 6 s. These slow-moving ejecta reflect convective motion near the BH.
@@
Figure 9: Time evolution of the unbound mass reaching $r=10^{3}$ km, as defined in Equation (1), for various models. Each model is represented by a different color, as indicated in the legend.
@@
We analyze tracers that are initially distributed with the fluid density. Following the gravitational binding criterion defined in Section II.3, Figure 9 shows the temporal evolution of the cumulative unbound mass reaching $r=10^{3}$ km for different models, as indicated in the legend. Models that include heating exhibit rapid unbound mass growth, reaching saturation within $\sim 0.5$ s. In contrast, non-heating (nH) models (light blue and light red lines) gradually increase over roughly 2 s. This points out that neglecting $r$-process heating results in slower ejection of material, with the saturation timescale differing by 1.5 s. The final unbound mass values ($M_{ej}$) are listed in Table 1, showing a 10% increase in models that include heating. Additionally, models with higher viscosity ($\alpha=0.06$) evolve more quickly and eject 10% more mass than those with $\alpha=0.03$, consistent with the expectation that enhanced angular momentum transport facilitates earlier mass ejection.
VI Comparison with time-dependent uniform heating rate
@@ VII DISCUSSION
Angular momentum transport in realistic merger disks is expected to occur via magnetohydrodynamic (MHD) turbulence driven by the magnetorotational instability (e.g., Balbus and Hawley 1998) instead of shear viscosity. Previous long-term 3D MHD simulations of BH accretion disks from NS mergers have shown that such turbulence combined with large-scale magnetic stresses can drive faster and more massive disk winds for strong poloidal fields, with typical average outflow velocities around 0.1 c instead of 0.03 c for viscous hydrodynamic models (Siegel and Metzger, 2018; Fernández et al., 2019; Miller et al., 2019; Just et al., 2022; Kiuchi et al., 2023). In this regime of higher baseline velocities, the relative impact of r-process heating is likely to be reduced, leading to more moderate velocity increases of the order of $\sim 10\%$, if the energy gain per baryon remains constant, compared to the factor of two enhancement observed in our simulations. For toroidal fields or weak poloidal fields, average outflow velocities are closer to the viscous hydrodynamic values (Christie et al., 2019; Fahlman and Fernández, 2022; Hayashi et al., 2023), in which case $r$-process heating can still be an order unity correction to the outflow velocities. The field geometry immediately after the merger is complex, with a mixture of poloidal and toroidal components spanning a range of spatial scales due to turbulence (e.g., Most et al. 2021; Aguilera-Miret et al. 2023; Izquierdo et al. 2024; Kiuchi et al. 2024), hence the definitive impact of r-process heating on outflow velocities with realistic field geometries remains an area of active study.
Our current heating prescription interpolates the nuclear energy release based on $Y_{e,T6}$ and the instantaneous temperature. However, outflows from merger disks exhibit a broad range of entropies and expansion timescales, which also affect nucleosynthesis and heating rates. Within this framework, we expect that variations in the adopted heating parameters may lead to changes in the total nuclear energy release within a factor of three. Expanding the heating tables to include these additional parameters in future work can improve the physical accuracy of the heating rate.